OpenStudy (anonymous):

Use the table of the function f(x) = x4 - 2x3 to answer this question:

OpenStudy (anonymous):

x f(x) -1 3 0 0 1 -1 2 0 3 27 What is the average rate of CHANGE from x = -1 to x = 1? -2 0 2 4 @jim_thompson5910 Can you plz help me out!?

jimthompson5910 (jim_thompson5910):

what is f(-1) equal to?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

look for -1 in the x column

jimthompson5910 (jim_thompson5910):

then look to the right of it to find f(-1)

OpenStudy (anonymous):

oh sorry 3

jimthompson5910 (jim_thompson5910):

f(-1) means "f(x) when x = -1"

jimthompson5910 (jim_thompson5910):

yes f(-1) = 3

jimthompson5910 (jim_thompson5910):

f(1) = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

now compute \[\Large \frac{f(-1)-f(1)}{-1-1}\]

jimthompson5910 (jim_thompson5910):

I'm using the formula \[\Large \frac{f(a)-f(b)}{a-b}\] in this case, a = -1 and b = 1

jimthompson5910 (jim_thompson5910):

what do you get when you compute that fraction?

OpenStudy (anonymous):

idk what to do..

jimthompson5910 (jim_thompson5910):

\[\Large \frac{f(-1)-f(1)}{-1-1} = \frac{3-(-1)}{-1-1} = ??\]

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

-2? lol

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes you misplaced a negative sign somewhere

jimthompson5910 (jim_thompson5910):

the answer is -2

OpenStudy (anonymous):

yeah ahah thank you so much i feel dumb for doing that xD

jimthompson5910 (jim_thompson5910):

its ok, you're welcome